Problem 8-7 Develop a production plan and calculate the annual cost for a firm whose demand forecast is fall, 10,700; winter, 8,300; spring, 6,900; summer, 12,700. Inventory at the beginning of fall is 535 units. At the beginning of fall you currently have 35 workers, but you plan to hire temporary workers at the beginning of summer and lay them off at the end of summer. In addition, you have negotiated with the union an option to use the regular workforce on overtime during winter or spring only if overtime is necessary to prevent stockouts at the end of those quarters. Overtime is not available during the fall. Relevant costs are hiring, $110 for each temp; layoff, $220 for each worker laid off, inventory holding, $5 per unit-quarter; backorder, $10 per unit; straight time, $5 per hour; overtime, $8 per hour. Assume that the productivity is 0.5 unit per worker hour, with eight hours per day and 60 days per season. In each quarter, produce to the full output of your regular workforce, even if that results in excess production. In Winter and Spring, use overtime only if needed to meet the production required in that quarter. Do not use overtime to build excess inventory in prior seasons expressly for the purpose of reducing the number of temp workers in Summer. (Leave no cells blank - be certain to enter "0" wherever required. Negative values should be indicated by a minus sign. Round up "Number of temp workers, Workers hired and Workers laid off" to the next whole number and all other answers to the nearest whole number.) Fall. Winter Spring 6 900 Summer 12.700 Forecast 10.700 8.300
Fall 10,700 Winter Spring 6,900 Summer Forecast 8,300 -1765 12,700 535 0 1500 Beginning inventory 10165 6900 7765 12165 Production required 20330 15530 13800 24330 Production hours required 16800 16800 16800 16800 Production hours available1 0 10 Overtime hours 0 0 Temp workers C Temp worker hours available 0 16800 16800 Total hours available 16800 16800 10165 7765 Actual production 6365 12165 -1765 Ending inventory 0 1500 2500 Workers hired 0 0 Workers laid off Fall Winter Spring Summer Straight time $ Overtime
erv cot Fall Winter Spring Summer Straight time per Overtime Inventory Backorder Hiring Layoff Total $ m o Annual cost vic cal
Answer #1

A C H K L 1 Fall Winter Spring 6,900 Summer S Solver Parameters X 2 Forecast 10,700 8,300 12,700 3 Beginning inventory Produc

FORMULAS:

Fall Winter Spring Summer
Forecast 10700 8300 6900 12700
Beginning inventory 535 =C12 =D12 =E12
Production required =C2-C3 =D2-D3 =E2-E3 =F2-F3
Production hours required =C4/0.5 =D4/0.5 =E4/0.5 =F4/0.5
Production hours available =35*60*8 =35*60*8 =35*60*8 =35*60*8
Overtime hours 3330 3.01167188609594E-20
Temp workers =B8+C13-B14 =C8+D13-C14 =D8+E13-D14 =E8+F13-E14
Temp worker hours avl. =C8*60*8 =D8*60*8 =E8*60*8 =F8*60*8
Total hours available =C6+C7+C9 =D6+D7+D9 =E6+E7+E9 =F6+F7+F9
Actual production =C10*0.5 =D10*0.5 =E10*0.5 =F10*0.5
Ending inventory =C3+C11-C2 =D3+D11-D2 =E3+E11-E2 =F3+F11-F2
Workers hired 0 12
Workers laid off 0 =F13
Costs
Straight time 5 =(C6+C9)*$B17 =(D6+D9)*$B17 =(E6+E9)*$B17 =(F6+F9)*$B17
Overtime 8 =$B18*C7 =$B18*D7 =$B18*E7 =$B18*F7
Inventory 5 =$B19*MAX(0,C12) =$B19*MAX(0,D12) =$B19*MAX(0,E12) =$B19*MAX(0,F12)
Backorder 10 =$B20*MAX(0,-C12) =$B20*MAX(0,-D12) =$B20*MAX(0,-E12) =$B20*MAX(0,-F12)
Hiring 110 =$B21*C13 =$B21*D13 =$B21*E13 =$B21*F13
Layoff 220 =$B22*C14 =$B22*D14 =$B22*E14 =$B22*F14
Total =SUM(C17:C22) =SUM(D17:D22) =SUM(E17:E22) =SUM(F17:F22)
Annual Cost = =SUM(C23:F23)